How do you solve a little tiny, micro, algebreic equation if it just happen to be 3X+34+3X+4-2X+5X-23-+34=1243
Answer by stanbon(75887)

3X+34+3X+4-2X+5X-23-+34=1243
Combine "like" terms to get:
9x -19 = 1243
Add 19 to each side to get:
9x = 1262
Divide both sides by 9 to get:
x = 1262/9
====================
Cheers,
Stan H.